OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later. A. 53 B. 63 C. 61 D. 59

OpenStudy (anonymous):

53

OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5. A. y = 6x + 5 B. y = -5/6 x + 5 C. y = -6x + 5 D. y = 5/6 x + 5

OpenStudy (anonymous):

B